Names and Addresses

The symbol ⊲ means “participant”;^{⎕} means “position paper (co)author”.

Birger Andersen Department of Computer Science University of Copenhagen Universitetsparken 1 2100 Copenhagen Denmark

Status: ⊲ ^{⎕}
Tel:  +45 31 39 64 66
Fax:  +45 31 39 02 21
E-mail: birger@diku.dk

Andrew P. Black Digital Equipment Corporation Cambridge Research Laboratory 1 Kendall Square, Building 700 Cambridge, Massachusetts 02139 USA

Status: ⊲ ^{⎕}
Tel:  +1 617 621 6633
Fax:  +1 617 621 6650
E-mail: black@crl.dec.com

Gregor v. Bochmann Departement d’Informatique et de Recherche Operationnelle Universite de Montreal Canada

Status: ⊲ ^{⎕}
Tel:  +1 514 343 7484
Fax:  +1 514 343 5834
E-mail: bochmann@iro.umontreal.ca

Gilad Bracha Computer Science Department University of Utah Salt Lake City, UT 84112 USA

Status: ⊲
Tel:  +1 801 521 5211

Fax:

E-mail: gilad@cs.utah.edu

Simon Brock School of Information Systems UEA Norwich NR4 7TJ England

Status: ⊲
Tel:  +44 603 593 164

Fax:

E-mail: shb@sys.uea.ac.uk

Brian Brown Methodologie du Logiciel Bull SA F6/0D/07 Rue Jean-Jaures F78340 les Clayes-sous-Bois France

Status: ⊲
Tel:  3080-6382
Fax:  3080-7078
E-mail: brian.brown@frcl.bull.fr

Luca Cardelli Digital Equipment Corporation Systems Research Center 130 Lytton Avenue Palo Alto CA 94301 USA

Status: ^{⎕}
Tel:  +1 415 853 2100
Fax:  +1 415 324 4873
E-mail: luca@src.dec.com

David Carrington Software Verification Research Centre Department of Computer Science The University of Queensland Australia

Status: ⊲ ^{⎕}

Tel:

Fax:  +61 7 365 1999
E-mail: davec@cs.uq.oz.au

Bruce A. Conrad WordPerfect Corporation 1555 North Technology Way Orem, Utah 84057 USA

Status: ⊲ ^{⎕}
Tel:  +1 801 228 7006
Fax:  +1 801 222 5077
E-mail: conrad@bert.cs.byu.edu

Rainer Fischbach Hirtenbrünnle 25 D-7245 Starzach Germany

Status: ⊲ ^{⎕}
Tel:  +49 7478 2149
Fax:  +49 7478 2148
E-mail: rf@garfield.t-informatik.ba-stuttgart.de

Andreas V. Hense Fachbereich 14 Universitaet des Saarlandes Im Stadtwald D-6600 Saarbruecken 11 Germany

Status: ⊲ ^{⎕}
Tel:  +49 681 302 2464

Fax:

E-mail: hense@cs.uni-sb.de

John S. Hogg Mail stop QE 012 Bell-Northern Research Ltd. P.O. Box 3511, Station C Ottawa, Ontario K1Y 4H7 Canada

Status: ⊲
Tel:  +1 613 763 7197
Fax:  +1 613 763 5647
E-mail: hogg@bnr.ca

Urs Hölzle Stanford University Center for Integrated Systems RM 042 Stanford, CA 94305-4070 USA

Status: ⊲ ^{⎕}
Tel:  +1 415 725 3695
Fax:  +1 415 725 6949
E-mail: hoelzle@cs.stanford.edu

Norman C. Hutchinson Department of Computer Science University of British Columbia 6356 Agricultural Road Vancouver, B.C. V6T 1Z2 Canada

Status: ⊲ ^{⎕}
Tel:  +1 604 224 8188

Fax:

E-mail: norm@cs.ubc.ca

Eric Jul DIKU Universitetsparken 1 2100 Kbh Denmark

Status: ⊲
Tel:  +45 3139 6466
Fax:  +45 3139 0221
E-mail: eric@diku.dk

Christian Laasch Department of Computer Science Information Systems - Databases ETH-Zürich CH-8092 Zürich Switzerland

Status: ⊲ ^{⎕}
Tel:  +41 1 254 7243
Fax:  +41 1 262 3973
E-mail: laasch@inf.ethz.ch

Serge Lacourte Bull-IMAG Systèmes 2 rue de Vignate 38610 Gières France

Status: ⊲
Tel:  +33 7654 4912
Fax:  +33 7654 7615
E-mail: lacourte@mururoa.imag.fr

Doug Lea SUNY at Oswego NY CASE Center NY 13126 USA

Status: ⊲ ^{⎕}
Tel:  +1 315 341 2688

Fax:

E-mail: dl@oswego.edu

Emmanuel Lenormand Bull-IMAG Systèmes 2 rue de Vignate 38610 Gières France

Status: ^{⎕}
Tel:  +33 7654 4912
Fax:  +33 7654 7615
E-mail: lenorman@mururoa.imag.fr

Karl Lieberherr Northeastern University College of Computer Science 360 Huntington Avenue Boston, MA 02115 USA

Status: ⊲ ^{⎕}
Tel:  +1 617 437 2077
Fax:  +1 617 437 5121
E-mail: lieber@corwin.ccs.northeastern.edu

Ole Lehrmann Madsen Computer Science Department Aarhus University Ny Munkegade DK–8000 Aarhus C Denmark

Status: ⊲ ^{⎕}
Tel:  +45 8612 7188
Fax:  +45 8613 5725
E-mail: olmadsen@daimi.aau.dk

Boris Magnusson Department of Computer Science University of Lund PO Box 118, S–221 00 Lund Sweden

Status: ⊲ ^{⎕}
Tel:  +46 4610 8044
Fax:  +46 4613 1021
E-mail: boris@dna.lth.se

Gianni Mainetto CNUCE - Institute of CNR Via S. Maria, 36 56126 Pisa Italy

Status: ⊲
Tel:  +39-50-593285
Fax:  +39-50-576751
E-mail: gmsys@icnucevm.cnuce.cnr.it

W. B. (Rick) Mugridge Department of Computer Science University of Auckland Auckland New Zealand

Status: ⊲ ^{⎕}
Tel:  NZ+9 737 999 x8914
Fax:  NZ+9 737 934
E-mail: rick@cs.aukuni.ac.nz

Birger Møller-Pedersen Norwegian Computing Center Gaustadalleen 23 P.O.Box 114 Blindern N–0314 Oslo 3 Norway

Status: ⊲ ^{⎕}
Tel:  +47 2 45 3500
Fax:  +47 2 69 7660
E-mail: birger@nr.no

Jens Palsberg Computer Science Department Aarhus University Ny Munkegade DK–8000 Aarhus C Denmark

Status: ⊲ ^{⎕}
Tel:  +45 8612 7188
Fax:  +45 8613 5725
E-mail: palsberg@daimi.aau.dk

Oskar Permvall Department of Computer Science Lund University PO Box 118, S–221 00 Lund Sweden

Status: ⊲ ^{⎕}
Tel:  +46 4610 8044
Fax:  +46 4613 1021
E-mail: oskar.permvall@dna.lth.se

Michel Riveill Bull-IMAG Systèmes 2 rue de Vignate 38610 Gières France

Status: ^{⎕}
Tel:  +33 7654 4912
Fax:  +33 7654 7615
E-mail: riveill@mururoa.imag.fr

Markku Sakkinen Department of Computer Science And Information Systems University of Jyväskylä PL 35, SF-40351 Jyväskylä Finland

Status: ⊲ ^{⎕}
Tel:  +35 841 603 016

Fax:

E-mail: sakkinen@jytko.jyu.fi

Marc H. Scholl Department of Computer Science Information Systems - Databases ETH-Zürich CH-8092 Zürich Switzerland

Status: ^{⎕}
Tel:  +41 1 254 7243
Fax:  +41 1 262 3973
E-mail: scholl@inf.ethz.ch

Michael I. Schwartzbach Computer Science Department Aarhus University Ny Munkegade DK–8000 Aarhus C Denmark

Status: ⊲ ^{⎕}
Tel:  +45 8612 7188
Fax:  +45 8613 5725
E-mail: mis@daimi.aau.dk

Alan Snyder Hewlett-Packard Laboratories P.O. Box 10490 Palo Alto CA 94306-0971 USA

Status: ⊲
Tel:  +1 415 857 8764
Fax:  +1 415 857 8526
E-mail: snyder@hpl.hp.com

Clemens A. Szyperski Institut für Computersysteme ETH-Zentrum, CH-8092 Zürich Switzerland

Status: ⊲
Tel:  +41 1 254 7318
Fax:  +41 1 262 3973
E-mail: szyperski@inf.ethz.ch

Andrew Watson APM Ltd Poseidon House Castle Park Cambridge CB3 0RD UK

Status: ⊲ ^{⎕}
Tel:  +44 223 323010
Fax:  +44 223 359779
E-mail: ajw@ansa.co.uk

Alan Wills Department of Computer Science The University of Manchester Manchester M13 9PL United Kingdom

Status: ⊲ ^{⎕}
Tel:  +44 61 275 6135

Fax:

E-mail: alan@cs.man.ac.uk